11:00 Eastern; 08:00 Project/Pacific Time. bluejeans.com/426716450

Attendees

Notes

Merlin Fisher-Levine :

  • Lots of hours, not a lot of ticking off milestones.
  • Away Thursday & Friday last week, Friday the week before.
  • Met with Roodman, Charles, et al from the Camera Team (DM-22009).
    • Worried about what has been happening in terms of cleaning, etc.
    • Deliverable to compare defects that Camera Team find and that DM find.
    • And to compare that the master biases generated by DM are the same as those found by the Camera Team.
  • Discussion of what BOT data it's useful for DM to reprocess, per message from Robert Lupton.
    • Seems unlikely that we want to repeat all Camera analysis on all data.
    • Will pin down Robert Luptonto get a “punch list” written, then try to get him on a call.
  • Started on reviewing Craig's PR (DM-18683); should be done soon.
  • Then starting on Spectrator.
    • Take “real data” files that Jeremy has processed and compare with results of running our Spectractor task on the same data.
    • Should correctly pull calspec standards, etc, from Simbad.
    • Needs to be ticketed.

Andrés Alejandro Plazas Malagón :

  • Meeting with Camera Team, as above.
    • Has some action points on DM-22009 which he would like to do this week.
    • This will be a parent ticket, for some subsidiary tickets which Andrés will file this week.
  • DM-19903 — columnar defects.
    • Addressing review comments.
    • Ongoing discussion on #dm-science-pipelines this morning, but there is general agreement that this isn't a major point.
    • Will have to address the problem identified by Tim J, but elegance/efficiency is not vital.
    • Will call John Swinbankif stuck.
  • Gen 3 conversion of PTC task stuck due to DM-21981.

John Swinbank :

  • Away next week; back on 18 November. Next meeting then.